OpenStudy (camijacobson):

What is the slope of the line that passed through the points (-2,1) and (1,6)? A. 3/5 B. -5 C. 5/3 D. -7

OpenStudy (black_ninja123):

Do you know the slope formula?

OpenStudy (black_ninja123):

\[\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (black_ninja123):

x1=-2 y1=1 x2=1 y2=6

OpenStudy (black_ninja123):

\[\frac{ 6-1 }{ 1-(-2) }\]

OpenStudy (black_ninja123):

\[\frac{ 6-1 }{ 1+2 }\]

OpenStudy (black_ninja123):

\[\frac{ 5 }{ 3 }\]